OpenStudy (anonymous):

differential equations....use the method of variation of parameters to find a particular solution of the given equation y'' - 2y' + y = 3e^(3x) differential equations....use the method of variation of parameters to find a particular solution of the given equation y'' - 2y' + y = 3e^(3x) @Mathematics

OpenStudy (anonymous):

i get C1 e^x + c2e^x x + e^(3x).....that sound right to anyone?

OpenStudy (jamesj):

That's correct.

OpenStudy (anonymous):

woot woot!

OpenStudy (jamesj):

1 is a repeat root, so c1.e^x + c2.x.e^x are the homogeneous solutions

OpenStudy (anonymous):

only took me an hour to get that lol

OpenStudy (amistre64):

i get to take a real class on ODEs next term yay!!

OpenStudy (anonymous):

it's kinda interesting...difficult concepts but interesting none the less

OpenStudy (amistre64):

there is one that i can almost recall, v=y/x ... but what is that called?

OpenStudy (anonymous):

homogeneous equation

OpenStudy (anonymous):

james...i wrote the equation wrong

OpenStudy (anonymous):

it's y'' - 2y' + y = 4e^(3x)

OpenStudy (jamesj):

So yes if y = e^3x, then we have y'' - 2y' + y = (9-6+1)e^3x = 4e^3x

OpenStudy (jamesj):

Ok, so yes indeed y = c1.e^x + c2.x.e^x + e^3x is the general solution.

OpenStudy (anonymous):

thank....god

OpenStudy (anonymous):

i got 4 left...i gotta work on them...u gonna be online so u can double check me by chance?

OpenStudy (jamesj):

Probably, yes.

OpenStudy (anonymous):

great...thx for your help